Rogelio and Jin are looking up at the top of the tower that is 36.6m tall. If Jin is looking up at an angle of elevation of 24 degree, and Rogelio is on the opposite side of the tower looking up at an angle of 17 degree, How far apart are Rogelio and Jin to the nearest meter? Assume their eyes are both 1.8m meter off the ground.
Answer by Cromlix(4381)

Consider the problem as two right angled
triangles back to back.
Rogelio stands on one side of the tower
and Jin stands on the other.
Now we have the distance from Rogelio
to the tower and the distance of Jin
from the tower. These represent the
adjacent side of both the triangles.
The opposite side is the tower, but as
both men's eyes are 1.8 metres from the ground
then the opposite side is given the value
36.6 - 1.8 = 34.8 metres.
We will use the tan ratio as we have an angle
and an opposite side. We must find the lengths
of the two adjacents.
Rogelio:
Tan 17 = 34.8/adjacent
Adjacent = 34.8/tan 17 = 113.8 metres
Jin
Tan 24 = 34.8/adjacent
Adjacent = 34.8/tan 24 = 78.2 metres
Add the two adjacent values
113.8 + 78.2 = 192 metres.
This is how far apart Rogelio and Jin are.
Hope this helps:-)